The length of a rectangular garden is 2 feet longer than 3 times its width. If the perimeter of the garden is 100 feet, find the width and length of the garden
Answer by ewatrrr(24785)

 
Hi

length of a rectangular garden is 2 feet longer than 3 times its width
Let x and (3x+2) represent the width and length respectively:
L + W = P/2 = 100/2 = 50
(3x+2) + x = 50
4x = 48
x = 12ft, the width, the length is 38ft